9. [7 points] Use a Direct Comparison Test to determine if the series converges or diverges. 00 7n 4n - 2 7121
Answer #1

I am attaching the pictures of solution. If you find my answer helpful please press thumbs up button. Thanks

Ans Direct Comparison Test : Suppose two series Zan and E by are such that ang bnyo for alln bu san for all n. and if Then Eч 7) ЛЈою 70 20 м) — 3 and чи м7) 2) –72 20 17 ) 2) ч 4 7° 2% м7 ) 7o 263 2 1 3) &0) I because so 2. у } hal 2 a70, 270Now (4u zal uh > Enh (и то -) — u uh ч- 2 4 ikust uh а) 7 7 ч Тесалле ->0 24 м, ) and if bcc and - ab cac azO а) bn Lan (211 SW IT n-1 O - Å bn n-1 کے n=1 m=) -) HD it is geometric series of the form & arnt which is convergent for lrk , divergentNow 141= 1 / 1 => 1 = I> so E bn n=1 (515) / is divergent Now bay 6, 6.0 bny, o and an7o by <an for nyl Also bn is divergent.given series is divergent